OpenStudy (anonymous):

Simplify x^2+8x/x^2: A. x+8/x^2 B. x+8/x C. 8x D. 8

OpenStudy (anonymous):

if you factor an x from the top and bottom you'll cancel an x... another way is to split the numerator \[\frac{x^2}{x^2}+\frac{8x}{x^2}\] = \[1+\frac{8}{x}\]

OpenStudy (anonymous):

if you put them under the same denominator you have to multiply the 1 by x which gets you back to your answers \[\frac{x+8}{x}\]

OpenStudy (anonymous):

why they don't teach splitting the numerators when doing this baffles me because you use it in calculus a lot

OpenStudy (anonymous):

Well, I guess I'm getting a tip for Calculus in 2 years. Thanks!